OpenStudy (anonymous):

A salt is known to be an alkali metal fluoride. A quick approximate freezing point determination indicates that 4 g of salt dissolved in 100 g of water produces a solution which freezes at about -2.8 degrees celsius. What is the formula of the salt? (assume i=2)

OpenStudy (anonymous):

can someone please do this for me i tried working it out but i can't figure it out

OpenStudy (anonymous):

did you use dT = Kf * m * i first you get out molality m = dT/Kf*i and then amount of solute is m = n(solute) / m(water) n (solute) = m(molality) * m(water) and then n= m/M -> M = m/n